Calculate the overall complex dissociation equilibrium constant for the Cu(NH3)42+ ion, given that ß4 for this complex is 2.1 × 1013

Grade:9

2 Answers

Sunil Kumar FP
9 years ago
the overall complex dissociation equilibrium constant for the Cu(NH3)42+ ion
the overall dissociation constant is the reciprocal of overall stability constant i,e 1/beta
=1/2.1*10^13 =4.7 *10^-14
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
The overall complex dissociation equilibrium constant for the Cu(NH3)42+ ion
The overall dissociation constant is the reciprocal of overall stability constant i,e 1/beta =1/2.1 x 1013 =4.7 x 10-14
 
Thanks and regards,
Kushagra